# SAMG: State-Action-Aware Offline-to-Online Reinforcement Learning with Offline Model Guidance

### v1

@@ Introduction
Ablation study of distribution-aware weight
@@
The state-action-aware weight $\alpha(s,a)$ provides an estimate of the offline degree given a $(s,a)$ pair and is instantiated as $p_{o}$ of the C-VAE model. To demonstrate the impact of the state-action-aware weight, we compare several different architectures on the environment Halfcheetah-medium with CQL-SAMG, including: (i) SAMG, (ii) 1/2 for online and 1/2 for offline (denoted as SAMG-even), (iii) random probability (generate random probability, denoted as SAMG-random), (iv) the vanilla RL algorithms.
The casual selection of C-VAE decreases the algorithm performance during the initial training phase which is even notably inferior to the vanilla algorithms and can never exceed the SAMG performance (Figure 5). Moreover, SAMG with casual C-VAE catches up with and surpasses the performance of the vanilla algorithms, which highlights the advantage of SAMG paradigm. SAMG manages to eliminate the influence of offline datasets, so once the algorithm adapts to the new environment, the algorithm is able to explore the environment and achieve higher scores. SAMG improves over the SAMG-even and vanilla algorithms by 15.3% and 5.2%, respectively. Furthermore, the state-action-aware weight represents a class of functions, as stated in Section SAMG: methodology. Given that $\alpha$ represents a class of function, we provide some other possible implementation techniques, as detailed in the Appendix A.2.
Sensitivity analysis of threshold $p_{om}$
It is a crucial hyperparameter that $p_{om}$ holds the lower threshold of OOD samples. We evaluate the sensitivity of $p_{om}$ on environment Antmaze-large-diverse with IQL-SAMG across a range of numbers, including 0, 0.3, 0.5, 0.75 (normal value). The results are shown in Figure 5. We find that $p_{om}$ is more sensitive in more complex environments, like Antmaze. Inappropriate selection of $p_{om}$ may lead to a decline in algorithm performance, or even complete divergence, thus highlighting the importance of $p_{om}$.
Does SAMG rely on offline data coverage?
@@
This part aims to evaluate the relationship between algorithm performance improvement and the coverage rate of the offline dataset to showcase the effect. To model the data coverage of the offline dataset, We apply t-SNE (Van der Maaten and Hinton 2008) to perform dimensionality reduction, then cluster with t-SNE information on all levels of datasets of a given environment, as detailed in Appendix C.7.
@@
The results illustrate that SAMG indeed correlates with the sample coverage rate (See Figure 5, where The left represents the medium-replay level, the middle represents the medium level and the right represents the medium-expert level of the dataset). Middle sample coverage rate yields greater performance improvement for SAMG, while higher or lower coverage rates lead to less significant performance improvements. This is because that lower coverage rates induce a narrow distribution of the offline dataset, resulting in limited information and accuracy of the offline model. Conversely, higher coverage rates often contribute to satisfaction with the offline model, hence further improvement is limited. However, SAMG consistently achieves improvements over vanilla algorithms and moderate sample coverage rates are common in practical applications, highlighting the superiority of SAMG.
Conclusion, limitation and future work
@@
In this paper, a novel paradigm named SAMG is proposed to eliminate the tedious usage of offline data and leverage the pre-trained offline model instead, thereby ensuring 100% online sample utilization and better fine-tuning performance. SAMG freezes the offline critic and combines online and offline critics with a state-action-aware coefficient. The state-action-aware coefficient models the complex distribution of the offline dataset and provides the probability of a given state-action sample deviating from the offline distribution. Theoretical analysis proves the convergence optimality and lower estimation error. Experimental results demonstrate the superiority of SAMG over vanilla baselines. However, it can be noticed that performance improvement of SAMG is limited if the offline dataset distribution is extremely narrow. This limitation could potentially be mitigated by designing specific update strategies for OOD samples, which serves as an interesting direction for future work.
Appendices
Appendix A A. Theoretical analysis
A.1. Contraction property
@@
Our algorithm actually breaks the typical Bellman Equation of the RL algorithm denoted as $Q=\max\limits_{\pi\in\Pi}\mathcal{B}^{\pi}=\max\limits_{\pi\in\Pi}(r_{\pi}+\gamma P_{\pi}Q)$. Instead we promote Equation 6. In order to prove the convergence of the updating equation, we introduce the contraction mapping theorem which is widely used to prove the convergence optimality of RL algorithm.
Theorem 3 (Contraction mapping theorem).
@@
We just need to prove that this equation satisfies the contraction property of theorem  3 and naturally we can ensure the convergence of the algorithm.
Take the right hand of equation  (6) as function $f(Q)$ and consider any two vectors $Q_{1},Q_{2}\in\mathbb{R}^{\mathcal{S}}$, and suppose that $\pi_{1}^{*}\doteq\mathop{\arg\max}\limits_{\pi}(r_{\pi}+\gamma P_{\pi}Q_{1})$, $\pi_{2}^{*}\doteq\mathop{\arg\max}\limits_{\pi}(r_{\pi}+\gamma P_{\pi}Q_{2})$. Then,
@@
$\begin{split}&f(Q_{1})-f(Q_{2})\\ =&(1-p^{\pi_{1}^{*}})\mathcal{B}^{\pi_{1}^{*}}Q_{1}+p^{\pi_{1}^{*}}\mathcal{B}^{\pi_{1}^{*}}Q^{off}\\ &-\left[(1-p^{\pi_{2}^{*}})\mathcal{B}^{\pi_{2}^{*}}Q_{2}+p^{\pi_{2}^{*}}\mathcal{B}^{\pi_{2}^{*}}Q^{off}\right]\\ \leq&(1-p^{\pi_{1}^{*}})\mathcal{B}^{\pi_{1}^{*}}Q_{1}+p^{\pi_{1}^{*}}\mathcal{B}^{\pi_{1}^{*}}Q^{off}-\\ &\left[(1-p^{\pi_{1}^{*}})\mathcal{B}^{\pi_{1}^{*}}Q_{2}+p^{\pi_{1}^{*}}\mathcal{B}^{\pi_{1}^{*}}Q^{off}\right]\\ =&(1-p^{\pi_{1}^{*}})(\mathcal{B}^{\pi_{1}^{*}}Q_{1}-\mathcal{B}^{\pi_{1}^{*}}Q_{2})\\ =&\gamma(1-p^{\pi_{1}^{*}})P^{\pi_{1}^{*}}(Q_{1}-Q_{2})\\ \leq&\gamma P^{\pi_{1}^{*}}(Q_{1}-Q_{2}).\end{split}$ (11)
@@
A.2. Convergence optimality
We consider a tabular setting for simplicity. We first write down the iterative form of Equation 4 as below:
@@ A.2. Convergence optimality
$\begin{split}&Q_{k+1}(s,a)=Q_{k}(s,a)\\ &-\alpha_{k}(s,a)\left[Q_{k}(s,a)-\left(r_{k+1}+\gamma Q_{k}(s_{k+1},a_{k+1})\right)\right]\\ &-\alpha_{k}(s,a)\gamma p(s,a)\left(Q^{off}(s_{k+1},a_{k+1})-Q_{k}(s_{k+1},a_{k+1})\right).\end{split},$ (13)
@@
$\begin{split}\Delta_{k}(s,a)\doteq Q_{k}(s,a)-Q_{\pi}(s,a).\end{split}$ (15)
@@
$\begin{split}\Delta_{k+1}(s,a)=&\left(1-\alpha_{k}(s,a)\right)\Delta_{k}(s,a)\\ &+\alpha_{k}(s,a)\eta_{k}(s,a),\quad s=s_{k},a=a_{k}.\end{split}$ (16)
@@
$\begin{split}&\eta_{k}(s,a)\\ =&\left[r_{k+1}+\gamma Q_{k}(s_{k+1},a_{k+1})-Q_{\pi}(s,a)\right]\\ &+\gamma p(s,a)\left(Q^{off}(s_{k+1},a_{k+1})-Q_{k}(s_{k+1},a_{k+1})\right)\\ =&\left[r_{k+1}+\gamma Q_{k}(s_{k+1},a_{k+1})-Q_{\pi}(s,a)\right]\\ &+\gamma p(s,a)\left\{\left[Q^{off}(s_{k+1},a_{k+1})-Q_{\pi}(s_{k+1},a_{k+1})\right]\right.\\ &\left.+\left[Q_{\pi}(s_{k+1},a_{k+1})-Q_{k}(s_{k+1},a_{k+1})\right]\right\}\\ =&\underbrace{\left[r_{k+1}+\gamma Q_{k}(s_{k+1},a_{k+1})-Q_{\pi}(s,a)\right]}_{\Gamma_{1}}\\ &-\underbrace{\gamma p(s,a)\left(Q_{k}(s_{k+1},a_{k+1})-Q_{\pi}(s_{k+1},a_{k+1})\right)}_{\Gamma_{2}}\\ &+\underbrace{\gamma p(s,a)\left(Q^{off}(s_{k+1},a_{k+1})-Q_{\pi}(s_{k+1},a_{k+1})\right)}_{\Gamma_{3}}.\end{split}$ (17)
Similarly, deducting $Q_{\pi}(s,a)$ from both side of Equation 14 gets:
$\begin{split}\Delta_{k+1}(s,a)=&\left(1-\alpha_{k}(s,a)\right)\Delta_{k}(s,a)\\ &+\alpha_{k}(s,a)\eta_{k}(s,a),\quad s\neq s_{k}\ or\ a\neq a_{k}.\end{split}$
@@
Theorem 4 (Dvoretzky’s Throrem).
@@
$\begin{split}\mathbb{E}[\Gamma_{1}]=&\mathbb{E}\left[r_{k+1}+\gamma Q_{k}(s_{k+1},a_{k+1})-Q_{\pi}(s_{k},a_{k})\left|s_{k},a_{k}\right.\right]\\ =&\mathbb{E}\left[r_{k+1}+\gamma Q_{k}(s_{k+1},a_{k+1})\left|s_{k},a_{k}\right.\right]-Q_{\pi}(s_{k},a_{k}).\end{split}$
Since $Q_{\pi}(s_{k},a_{k})=\mathbb{E}\left[r_{k+1}+\gamma Q_{\pi}(s_{k+1},a_{k+1})\left|s_{k},a_{k}\right.\right]$, the above equation indicates that,
$\begin{split}\mathbb{E}[\Gamma_{1}]=&\gamma\mathbb{E}\left[Q_{k}(s_{k+1},a_{k+1})-Q_{\pi}(s_{k+1},a_{k+1})\left|s_{k},a_{k}\right.\right].\end{split}$
@@
$\begin{split}\mathbb{E}[\Gamma_{2}]=&\gamma p(s_{k},a_{k})\mathbb{E}\left[Q_{k}(s_{k+1},a_{k+1})\right.\\ &-\left.Q_{\pi}(s_{k+1},a_{k+1})\left|s_{k},a_{k}\right.\right].\end{split}$
@@
$\begin{split}\mathbb{E}[\Gamma_{1}]-\mathbb{E}[\Gamma_{2}]=&\gamma\left(1-p(s_{k},a_{k})\right)\mathbb{E}\left[Q_{k}(s_{k+1},a_{k+1})-\right.\\ &\left.Q_{\pi}(s_{k+1},a_{k+1})\left|s_{k},a_{k}\right.\right].\end{split}$
@@
$\begin{split}&\big{\|}\mathbb{E}[\Gamma_{1}]-\mathbb{E}[\Gamma_{2}]\big{\|}_{\infty}\\ =&\gamma\left(1-p(s_{k},a_{k})\right)\\ &\big{\|}\sum\limits_{a^{\prime}\in\mathcal{A}}\sum\limits_{s^{\prime}\in\mathcal{S}}t\left(s^{\prime},a^{\prime}\left|s_{k},a_{k}\right.\right)\left|Q_{k}(s^{\prime},a^{\prime})-Q_{\pi}(s^{\prime},a^{\prime})\right|\big{\|}_{\infty}\\ =&\gamma\left(1-p(s_{k},a_{k})\right)\max\limits_{s^{\prime}\in\mathcal{S},a^{\prime}\in\mathcal{A}}\\ &\bigg{\{}\sum\limits_{a^{\prime}\in\mathcal{A}}\sum\limits_{s^{\prime}\in\mathcal{S}}t\left(s^{\prime},a^{\prime}\left|s_{k},a_{k}\right.\right)\left|Q_{k}(s^{\prime},a^{\prime})-Q_{\pi}(s^{\prime},a^{\prime})\right|\bigg{\}}\\ \leq&\gamma\left(1-p(s_{k},a_{k})\right)\sum\limits_{a^{\prime}\in\mathcal{A}}\sum\limits_{s^{\prime}\in\mathcal{S}}t\left(s^{\prime},a^{\prime}\left|s_{k},a_{k}\right.\right)\\ &\max\limits_{s^{\prime}\in\mathcal{S},a^{\prime}\in\mathcal{A}}\left[\left|Q_{k}(s^{\prime},a^{\prime})-Q_{\pi}(s^{\prime},a^{\prime})\right|\right]\\ =&\gamma\left(1-p(s_{k},a_{k})\right)\max\limits_{s^{\prime}\in\mathcal{S},a^{\prime}\in\mathcal{A}}\left[\left|Q_{k}(s^{\prime},a^{\prime})-Q_{\pi}(s^{\prime},a^{\prime})\right|\right]\\ =&\gamma\left(1-p(s_{k},a_{k})\right)\big{\|}\Delta_{k}(s,a)\big{\|}_{\infty}.\end{split}$

### v2

@@ 5 Experimental Results
Baselines. (i) SAMG algorithms, the SAMG paradigm is constructed on a variety of state-of-the-art O2O RL algorithms, including CQL (Kumar et al., 2020) and AWAC (Nair et al., 2020), IQL (Kostrikov et al., 2022). (ii) O2O RL algorithms, we implement the aforementioned O2O RL algorithms (CQL, AWAC and IQL). We also implement SPOT (Wu et al., 2022) (iii) Hybrid RL, we implement sota hybrid-RL-based algorithms, including Cal_QL (Nakamoto et al., 2023) and EDIS (Liu et al., 2024). (iv) Behavior Cloning (BC), we implement Behavior cloning based algorithm TD3+BC (Fujimoto and Gu, 2021). All algorithms are implemented based on CORL library (Tarasov et al., 2024) with implementation details in Appendix C.1. To ensure a fair comparison, all algorithms are pre-trained offline for 1M iterations followed by 200k iterations of online fine-tuning.
Benchmark tasks. We evaluate SAMG and the baselines across multiple benchmark tasks: (1) The Mujuco locomotion tasks (Fu et al., 2020), including three different kinds of environments (HalfCheetah, Hopper, Walker2d) where robots are manipulated to complete various tasks on three different levels of datasets. (2) The AntMaze tasks that an “Ant” robot is controlled to explore and navigate to random goal locations in six levels of environments. (3) The Adroit tasks include Pen, Door, and Relocate environments. Details are stated in Appendix C.3.
5.1 Empirical Results
The normalized scores of the vanilla algorithms with and without SAMG integrated are shown in Table 1. SAMG consistently outperforms the vanilla algorithms in the majority of environments, illustrating the superiority of SAMG. SAMG converges significantly faster than vanilla algorithms and can achieve higher performance. Notably, SAMG achieves the best performance with the simpler algorithm AWAC, while delivering substantial improvements with other algorithms. The reason for this counter-intuitive phenomenon is discussed in Appendix C.4, which just illustrates the effectiveness of SAMG. We present the cumulative regrets on Antmaze in Appendix C.5, further demonstrating the outstanding online sample efficiency of SAMG.
Although SAMG performs well in most environments, it is still worthwhile to notice SAMG may occasionally behave unsatisfactory (e.g., IQL-SAMG on HalfCheetah-medium-expert task). We discuss in Appendix C.6 that this exception is caused by the unique property of the environment rather than the defect of SAMG. We notice that the AWAC algorithm performs poorly in the Antmaze environment, resulting in SAMG struggling to initiate. This is because AWAC is relatively simple and not competent for the complex task of Antmaze; it is an inherent limitation of AWAC, rather than an issue with SAMG.
5.2 Ablation Analysis of the Coefficient
@@
The state-action-conditional coefficient $p(s,a)$, instantiated as $p^{vae}(s,a)$, estimates the offline degree for a given $(s,a)$ pair and is adaptively updated during training. To demonstrate the impact of this adaptive state-action-conditional coefficient, we compare several different architectures on the environment HalfCheetah with CQL-SAMG, including: (i) adopted SAMG setting (denoted as SAMG), (ii) static state-action-conditional coefficient (which means the VAE model is fixed once pre-trained offline, denoted as SAMG-static) (iii) 1/2 for online and 1/2 for offline (denoted as SAMG-even), (iv) randomly generated probability (denoted as SAMG-random), (v) the vanilla RL algorithms (denoted as Vanilla). The results are illustrated in Fig. 5.
SAMG shows consistent and significant improvement compared to other settings. Casual selections of C-VAE (SAMG-even and SAMG-random) exhibit notably inferior algorithm performance during the initial training phase, demonstrating the effectiveness of state-action-conditional coefficient structure. However, they catch up with and surpass the performance of the vanilla algorithms, highlighting the advantage of SAMG paradigm and offline information. SAMG improves over the SAMG-static and SAMG-even algorithms by 15.3% and 21.8% respectively.
Furthermore, we provide another possible implementation technique $p^{emb}$ given that $p(s,a)$ represents a function class. Refer to Appendix B.2 for complete details.
5.3 Sensitivity Analysis of Probability Threshold
It is a crucial hyperparameter that $p^{vae}_{m}$ holds the lower threshold of OOD samples. We evaluate the sensitivity of $p^{vae}_{m}$ on environment Antmaze with IQL-SAMG across a range of numbers, including 0.2, 0.5, 0.6, 0.7 (chosen value), and 0.8. The results are shown in Fig. 5. A slight change in the value of $p^{vae}_{m}$ (corresponding to 0.6 and 0.8) leads to a minor decrease in algorithm performance, but the impact is not significant. If the value is too large, the algorithm reverts to the vanilla version; if it is too small, it introduces incorrect offline critic perceptions of OOD samples, resulting in a performance decline. Furthermore, when $p^{vae}_{m}$ is small enough, all samples are regarded as OOD samples and updated solely based on offline knowledge, causing a significant drop in algorithm performance, which is also identical across different $p^{vae}_{m}$, as evidenced by the overlapping curves for 0.2 and 0.5 in Fig. 5.
5.4 Further Comparisons with Hybrid RL Algorithms
We compare SAMG with Hybrid RL based algorithms in Table 1. A natural problem arises: would the algorithm performance improve if the hybrid RL setting were replaced with offline-guidance setting? To illustrate this problem, we modify the Cal_QL algorithm with SAMG setting. Implementation details can be found in Appendix C.7. The results, presented in Table 2, indicate that SAMG still outperforms Cal_QL, demonstrating its superiority. As for EDIS algorithm, it relies heavily on the offline dataset, making it impractical to adapt to offline-guidance setting.
Table 2: Algorithm performance of Cal_QL and SAMG. The algorithms performance of Cal_QL compared to SAMG integrated Cal_QL algorithms. The result is the average normalized score of 5 random seeds $\pm$ (standard deviation).
Algorithms Hopper-mr Hopper-m Hopper-me Half-mr Half-m Half-me Walk-mr Walk-m Walk-me
Cal_QL 80.9(38.2) 78.1(8.7) 109.1(0.2) 51.6(0.8) 63.2(2.5) 95.6(4.3) 97.1(2.5) 83.6(0.8) 110.7(0.4)
SAMG 101.6(0.6) 99.8(2.1) 111.7(0.6) 56.4(1.5) 65.1(1.0) 96.3(0.5) 101.2(0.1) 97.8(1.7) 112.3(0.8)
Ant-u Ant-ud Ant-mp Ant-md Ant-lp Ant-ld Pen-b Door-b Relocate-b
Cal_QL 96.8(0.4) 63.8(43.4) 93.4(3.6) 94.0(2.2) 78.8(5.8) 73.0(19.4) -0.03(4.10) -0.33(0.01) -0.31(0.03)
SAMG 99.0(1.0) 66.4(23.0) 91.6(2.4) 96.0(1.6) 79.8(2.4) 72.4(11.2) 5.01(12.01) 1.35(0.57) 3.89(0.34)
5.5 Does SAMG Rely on Offline Data Coverage?
@@
This part aims to showcase the relationship between algorithm performance improvement and the coverage rate of the offline dataset. To model the data coverage of a specific offline dataset, we apply t-SNE (Van der Maaten and Hinton, 2008) to perform dimensionality reduction, then cluster with t-SNE information on all levels of datasets of a given environment, as detailed in Appendix C.8.
@@
The results illustrate that SAMG correlates with the sample coverage rate (See Fig. 5, where the left represents the medium-replay level, the middle represents the medium level and the right represents the medium-expert level of the dataset). Middle sample coverage rate yields more significant performance improvement for SAMG, while higher or lower coverage rates lead to less significant performance improvements. This is because lower coverage rates induce a narrow distribution of the offline dataset, resulting in limited information of the offline model. Conversely, higher coverage rates contribute to satisfaction with the offline model, hence further improvement is limited. However, moderate sample coverage rates are common in practical applications.
6 Related Work
Online RL with assistance of dataset. Expert-level demonstrations are considered and combined with online data to accelerate the online training process (Hester et al., 2018; Li et al., 2023; Nair et al., 2018; Rajeswaran et al., 2018; Vecerik et al., 2017). In contrast, another line of work trains the RL agent from scratch and accelerates the process by sampling both offline and online data (Hester et al., 2018; Song et al., 2023; Ball et al., 2023; Zhou et al., 2024a). Auxiliary behavioral cloning losses with policy gradients are utilized to guide the updates and accelerate convergence (Kang et al., 2018; Zhu et al., 2018, 2019).
Offline-to-online RL. Large task-agnostic datasets are utilized to pre-train a unified offline model and feed for subsequent online RL tasks (Aytar et al., 2018; Lowrey et al., 2019; Fujimoto et al., 2019; Baker et al., 2022; Lifshitz et al., 2024; Yuan et al., 2024).
A different category of work is to train offline with provided dataset and then fine-tune online in the same environment, which could be further divided as model-based O2O RL and model-free O2O RL. Model-based O2O RL algorithms combined with prioritized sampling scheme (Mao et al., 2022), adaptive behavior regularization algorithm MOTO (Rafailov et al., 2023) or energy-guided diffusion sampling technique (Liu et al., 2024) are proposed to mitigate O2O distribution shift.
For the model-free O2O RL, some offline RL algorithms are directly applied for O2O setting (Nair et al., 2020; Kumar et al., 2020; Kostrikov et al., 2022). A series of Q-ensemble based algorithms are proposed, while combined with balanced experience replay (Lee et al., 2022), state-dependent balance coefficient (Wang et al., 2024), uncertainty quantification guidance (Guo et al., 2023), uncertainty penalty and smoothness regularization (Wen et al., 2024) and optimistic exploration (Zhao et al., 2024). Recently some work attempts to efficiently explore the environment to accelerate the fine-tuning process: O3F optimistically takes actions with higher expected Q-values (Mark et al., 2022), PEX introduces an extra policy to adaptively explore and learn (Zhang et al., 2023), OOO framework maintains an exploration policy to collect data and an exploitation policy to train on all data (Mark et al., 2024) and PTGOOD utilizes planning procedure to explore high-reward areas distant from offline distribution (McInroe et al., 2024). There are some other independent works: SPOT brings out a density-based regularization term to model the behavior policy (Wu et al., 2022), Td3-BC integrates behavioural cloning /constraint that decays over time (Beeson and Montana, 2022), Cal-QL calibrates the learned Q-values at reasonable scale with some reference policy (Nakamoto et al., 2023). OLLIE proposes the O2O imitation learning (Yue et al., 2024).
We observe that a cocurrent work WSRL explores to initialize the replay buffer without retaining offline data (Zhou et al., 2024b). Nevertheless, SAMG proposes a complete and effective solution, whereas WSRL only explores the warm-up of the initial replay buffer.
7 Conclusion
@@
This paper proposes a novel paradigm named SAMG to eliminate the tedious usage of offline data and leverage the pre-trained offline critic model instead, thereby ensuring 100% online sample utilization and better fine-tuning performance. SAMG seamlessly combines online and offline critics with a state-action-conditional coefficient without introducing undesirable or questionable intrinsic rewards. This coefficient estimates the complex distribution of the offline dataset and provides the probability of a given state-action sample. Theoretical analysis proves the convergence optimality and lower estimation error. Experimental results demonstrate the superiority of SAMG over vanilla baselines. However, the performance improvement is limited if the offline dataset distribution is extremely narrow. This limitation could potentially be mitigated by designing specific update strategies for OOD samples, which is an interesting direction for future work.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Theoretical analysis
A.1 Contraction property
@@
Our algorithm actually breaks the typical Bellman Equation of the RL algorithm denoted as $Q=\max\limits_{\pi\in\Pi}\mathcal{B}^{\pi}=\max\limits_{\pi\in\Pi}(r_{\pi}+\gamma P_{\pi}Q)$. Instead we promote Eq. (1). In order to prove the convergence of the updating equation, we introduce the contraction mapping theorem which is widely used to prove the convergence optimality of RL algorithm.
Theorem A.1 (Contraction mapping theorem).
@@
We just need to prove that this equation satisfies the contraction property of theorem  A.1 and naturally we can ensure the convergence of the algorithm.
Take the right hand of Eq. ((1)) as function $f(Q)$ and consider any two vectors $Q_{1},Q_{2}\in\mathbb{R}^{\mathcal{S}}$, and suppose that $\pi_{1}^{*}\doteq\mathop{\arg\max}\limits_{\pi}(r_{\pi}+\gamma P_{\pi}Q_{1})$, $\pi_{2}^{*}\doteq\mathop{\arg\max}\limits_{\pi}(r_{\pi}+\gamma P_{\pi}Q_{2})$. Then,
@@
$\begin{split}&f(Q_{1})-f(Q_{2})\\ =&(1-p^{\pi_{1}^{*}})\mathcal{B}^{\pi_{1}^{*}}Q_{1}+p^{\pi_{1}^{*}}\mathcal{B}^{\pi_{1}^{*}}Q^{off}-\left[(1-p^{\pi_{2}^{*}})\mathcal{B}^{\pi_{2}^{*}}Q_{2}+p^{\pi_{2}^{*}}\mathcal{B}^{\pi_{2}^{*}}Q^{off}\right]\\ \leq&(1-p^{\pi_{1}^{*}})\mathcal{B}^{\pi_{1}^{*}}Q_{1}+p^{\pi_{1}^{*}}\mathcal{B}^{\pi_{1}^{*}}Q^{off}-\left[(1-p^{\pi_{1}^{*}})\mathcal{B}^{\pi_{1}^{*}}Q_{2}+p^{\pi_{1}^{*}}\mathcal{B}^{\pi_{1}^{*}}Q^{off}\right]\\ =&(1-p^{\pi_{1}^{*}})(\mathcal{B}^{\pi_{1}^{*}}Q_{1}-\mathcal{B}^{\pi_{1}^{*}}Q_{2})\\ =&\gamma(1-p^{\pi_{1}^{*}})P^{\pi_{1}^{*}}(Q_{1}-Q_{2})\\ \leq&\gamma P^{\pi_{1}^{*}}(Q_{1}-Q_{2}).\end{split}$ (11)
@@
A.2 Convergence optimality
We consider a tabular setting for simplicity. We first write down the iterative form of Eq. (5) as below:
@@ A.2 Convergence optimality
$\begin{split}Q_{k+1}(s,a)=&Q_{k}(s,a)-\alpha_{k}(s,a)\bigl{[}Q_{k}(s,a)-\left(r_{k+1}+\gamma Q_{k}(s_{k+1},a_{k+1})\right)\bigr{]}\\ &-\alpha_{k}(s,a)\gamma p(s,a)\bigl{(}Q^{off}(s_{k+1},a_{k+1})-Q_{k}(s_{k+1},a_{k+1})\bigr{)}.\end{split}$ (13)
@@
$\begin{split}\Delta_{k}(s,a)\doteq Q_{k}(s,a)-Q(s,\pi).\end{split}$ (15)
@@
$\begin{split}\Delta_{k+1}(s,a)=&\left(1-\alpha_{k}(s,a)\right)\Delta_{k}(s,a)+\alpha_{k}(s,a)\eta_{k}(s,a),\quad s=s_{k},a=a_{k}.\end{split}$ (16)
@@
$\begin{split}&\eta_{k}(s,a)\\ =&\bigl{[}r_{k+1}+\gamma Q_{k}(s_{k+1},a_{k+1})-Q^{\pi}(s,a)\bigr{]}+\gamma p(s,a)\bigl{[}Q^{off}(s_{k+1},a_{k+1})-Q_{k}(s_{k+1},a_{k+1})\bigr{]}\\ =&\bigl{[}r_{k+1}+\gamma Q_{k}(s_{k+1},a_{k+1})-Q^{\pi}(s,a)\bigr{]}+\\ &\gamma p(s,a)\Bigl{\{}\left[Q^{off}(s_{k+1},a_{k+1})-Q^{\pi}(s_{k+1},a_{k+1})\right]+\bigl{[}Q^{\pi}(s_{k+1},a_{k+1})-Q_{k}(s_{k+1},a_{k+1})\bigr{]}\Bigr{\}}\\ =&\underbrace{\bigl{[}r_{k+1}+\gamma Q_{k}(s_{k+1},a_{k+1})-Q^{\pi}(s,a)\bigr{]}}_{\Gamma_{1}}-\underbrace{\gamma p(s,a)\bigl{[}Q_{k}(s_{k+1},a_{k+1})-Q^{\pi}(s_{k+1},a_{k+1})\bigr{]}}_{\Gamma_{2}}\\ &+\underbrace{\gamma p(s,a)\bigl{[}Q^{off}(s_{k+1},a_{k+1})-Q^{\pi}(s_{k+1},a_{k+1})\bigr{]}}_{\Gamma_{3}}\\ =&\Gamma_{1}-\Gamma_{2}+\Gamma_{3}.\end{split}$ (17)
Similarly, deducting $Q^{\pi}(s,a)$ from both side of Eq. (14) gets:
$\begin{split}\Delta_{k+1}(s,a)=&\left(1-\alpha_{k}(s,a)\right)\Delta_{k}(s,a)+\alpha_{k}(s,a)\eta_{k}(s,a),\quad s\neq s_{k}\ or\ a\neq a_{k}.\end{split}$
@@
Theorem A.2 (Dvoretzky’s Throrem).
@@
$\begin{split}\mathbb{E}[\Gamma_{1}]=&\mathbb{E}\bigl{[}r_{k+1}+\gamma Q_{k}(s_{k+1},a_{k+1})-Q^{\pi}(s_{k},a_{k})\left|s_{k},a_{k}\right.\bigr{]}\\ =&\mathbb{E}\bigl{[}r_{k+1}+\gamma Q_{k}(s_{k+1},a_{k+1})\left|s_{k},a_{k}\right.\bigr{]}-Q^{\pi}(s_{k},a_{k}).\end{split}$
Since $Q_{\pi}(s_{k},a_{k})=\mathbb{E}\left[r_{k+1}+\gamma Q^{\pi}(s_{k+1},a_{k+1})\left|s_{k},a_{k}\right.\right]$, the above equation indicates that,
$\begin{split}\mathbb{E}[\Gamma_{1}]=&\gamma\mathbb{E}\bigl{[}Q_{k}(s_{k+1},a_{k+1})-Q^{\pi}(s_{k+1},a_{k+1})\left|s_{k},a_{k}\right.\bigr{]}.\end{split}$
@@
$\begin{split}\mathbb{E}[\Gamma_{2}]=&\gamma p(s_{k},a_{k})\mathbb{E}\bigl{[}Q_{k}(s_{k+1},a_{k+1})-Q^{\pi}(s_{k+1},a_{k+1})\left|s_{k},a_{k}\right.\bigr{]}.\end{split}$
@@
$\begin{split}\mathbb{E}[\Gamma_{1}]-\mathbb{E}[\Gamma_{2}]=&\gamma\left(1-p(s_{k},a_{k})\right)\mathbb{E}\bigl{[}Q_{k}(s_{k+1},a_{k+1})-Q^{\pi}(s_{k+1},a_{k+1})\left|s_{k},a_{k}\right.\bigr{]}.\end{split}$
@@
$\begin{split}&\big{\|}\mathbb{E}[\Gamma_{1}]-\mathbb{E}[\Gamma_{2}]\big{\|}_{\infty}\\ =&\gamma\left(1-p(s_{k},a_{k})\right)\Big{\|}\sum\limits_{a^{\prime}\in\mathcal{A}}\sum\limits_{s^{\prime}\in\mathcal{S}}t\left(s^{\prime},a^{\prime}\left|s_{k},a_{k}\right.\right)\bigl{|}Q_{k}(s^{\prime},a^{\prime})-Q^{\pi}(s^{\prime},a^{\prime})\bigr{|}\Big{\|}_{\infty}\\ =&\gamma\left(1-p(s_{k},a_{k})\right)\max\limits_{s^{\prime}\in\mathcal{S},a^{\prime}\in\mathcal{A}}\bigg{\{}\sum\limits_{a^{\prime}\in\mathcal{A}}\sum\limits_{s^{\prime}\in\mathcal{S}}t\left(s^{\prime},a^{\prime}\left|s_{k},a_{k}\right.\right)\bigl{|}Q_{k}(s^{\prime},a^{\prime})-Q^{\pi}(s^{\prime},a^{\prime})\bigr{|}\bigg{\}}\\ \leq&\gamma\left(1-p(s_{k},a_{k})\right)\sum\limits_{a^{\prime}\in\mathcal{A}}\sum\limits_{s^{\prime}\in\mathcal{S}}t\left(s^{\prime},a^{\prime}\left|s_{k},a_{k}\right.\right)\max\limits_{s^{\prime}\in\mathcal{S},a^{\prime}\in\mathcal{A}}\Bigl{[}\bigl{|}Q_{k}(s^{\prime},a^{\prime})-Q^{\pi}(s^{\prime},a^{\prime})\bigr{|}\Bigr{]}\\ =&\gamma\left(1-p(s_{k},a_{k})\right)\max\limits_{s^{\prime}\in\mathcal{S},a^{\prime}\in\mathcal{A}}\Bigl{[}\bigl{|}Q_{k}(s^{\prime},a^{\prime})-Q^{\pi}(s^{\prime},a^{\prime})\bigr{|}\Bigr{]}\\ =&\gamma\left(1-p(s_{k},a_{k})\right)\big{\|}\Delta_{k}(s,a)\big{\|}_{\infty}.\end{split}$